# Massive-STEPS: Massive Semantic Trajectories for Understanding POI Check-ins – Dataset and Benchmarks

### v2

@@
Abstract
Understanding human mobility through Point-of-Interest (POI) recommendation is increasingly important for applications such as urban planning, personalized services, and generative agent simulation. However, progress in this field is hindered by two key challenges: the over-reliance on older datasets from 2012-2013 and the lack of reproducible, city-level check-in datasets that reflect diverse global regions. To address these gaps, we present Massive-STEPS (Massive Semantic Trajectories for Understanding POI Check-ins), a large-scale, publicly available benchmark dataset built upon the Semantic Trails dataset and enriched with semantic POI metadata. Massive-STEPS spans 12 geographically and culturally diverse cities and features more recent (2017-2018) and longer-duration (24 months) check-in data than prior datasets. We benchmarked a wide range of POI recommendation models on Massive-STEPS using both supervised and zero-shot approaches, and evaluated their performance across multiple urban contexts. By releasing Massive-STEPS, we aim to facilitate reproducible and equitable research in human mobility and POI recommendation. The dataset and benchmarking code are available at: https://github.com/cruiseresearchgroup/Massive-STEPS.
1 Introduction
@@
Human mobility data and modeling are essential for understanding how individuals interact with and move through physical spaces. This understanding enables a wide range of applications, including urban planning [63], travel service recommendations [7], improved commercial advertising strategies [59], and Point-of-Interest (POI) recommendation [6, 22, 68]. Recently, human mobility data has become even more crucial with the increasing use of large language models (LLM) agents to simulate human-like behavior and routines [71, 20]. However, while simulated human mobility data are starting to gain popularity [31, 9], they may not accurately reflect real-world human behavior [33], highlighting the value of evaluating on real-world data. These advancements are enabled by and large with the rise of Location-based Social Networks (LBSNs), which generate vast amounts of spatio-temporal data through user check-ins  [68, 22]. This rich data source has allowed the development of POI recommendation systems that leverage users’ historical visiting behaviors to suggest relevant locations. Such systems enhance user engagement through personalization and provide commercial value to both users and businesses by aligning recommendations with individual preferences and available services [68, 6].
@@
Our paper addresses three critical gaps in POI recommendation research and datasets, as detailed in Section 2.1. First, as shown in Fig. 1, the field remains dominated by studies focused on just two cities, New York and Tokyo, based on the Foursquare dataset curated in [57]. This dataset, collected in 2012-2013, raises concerns about its temporal quality, as many POIs may no longer exist and user behavior may have changed [61]. While some recent studies have expanded to other cities [65, 28, 7], they often rely on the Global-scale Check-in Dataset (GSCD) [55, 56], which, despite its large coverage, is also from 2012-2013 and contains nearly 50% erroneous entries [29].
Second, most existing studies are difficult to reproduce, either due to the lack of clearly defined geographic boundaries or the unavailability of the datasets themselves, hindering fair comparison and replication. Finally, we join recent efforts [63] in advocating for the inclusion of low-resource and underrepresented cities in evaluation. Expanding beyond well-studied urban centers is essential for building more generalizable and universally applicable POI recommendation models. Table 1 summarizes these limitations in terms of geographic coverage, temporal span, and reproducibility.
Table 1: Comparison of check-in datasets commonly used for POI recommendation tasks. GSCD  [57, 56] and Semantic Trails [29] are global datasets not grouped into individual cities, whereas others perform city-level grouping. Replicable indicates whether city boundaries are clearly defined or can be reliably reconstructed.
@@
#cities Years #months POI Attributes Replicable Open-source
GSCD [57, 56] Varies 2012-2013 17    Coordinates, Category N/A
Semantic Trails [29] Varies 2012-2013, 2017-2018 24    Category N/A
NYC and Tokyo [57] 2 2012-2013 11    Coordinates, Category
Gowalla-CA [4, 62] 1 2009-2010 21    Coordinates, Category
AgentMove [7] 12 2012-2013 17    Coordinates, Category
Massive-STEPS 12 2012-2013, 2017-2018 24 Coordinates, Category, Name, Address
Massive-STEPS Dataset
In this paper, we introduced the Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) Dataset, derived from the Semantic Trails dataset (STD) [29]. Massive-STEPS includes high-quality check-ins from 2012-2013 and 2017-2018, providing more modern and updated POI check-in data. This supports longitudinal POI recommendation studies and addresses the limitations of older datasets commonly used in prior studies. The dataset covers 12 diverse cities across multiple regions, including East, West, and Southeast Asia, North and South America, Australia, the Middle East, and Europe. Notably, we placed a deliberate emphasis on under-explored regions by including cities such as Jakarta, Kuwait City, and Petaling Jaya, filling a key gap in POI recommendation research that has largely focused on major urban centers. We further enriched STD by aligning it with Foursquare’s Open Source Places dataset, incorporating metadata such as POI coordinates, POI names, and addresses—details unavailable in the original STD.
Benchmark Tasks
To demonstrate the utility of this dataset, we conducted an extensive benchmark on two tasks: (1) supervised POI recommendation and (2) zero-shot POI recommendation. Our benchmark covers a wide range of models, including traditional approaches, deep learning-based models, and more recent LLM-based methods. The goal of POI recommendation task is to predict a set of POIs that a user is likely to visit based on their current check-in trajectory and historical behavior. This reflects real-world applications such as personalized POI recommendations in location-based services. In addition, the scale of our dataset allows us to examine how urban features influence POI recommendation accuracy. Building on prior hypotheses, we propose a new insight: cities with more evenly distributed POI categories tend to be harder to model, as the absence of a dominant POI category makes user behavior less predictable.
Contribution
This paper introduces the Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) dataset, addressing gaps in existing POI recommendation research. Current POI check-in datasets are often only from 2012-2013, skewed to a few cities, and lack semantic metadata, hindering the development of robust and globally applicable models. While datasets like GSCD and STD offer broad geographic coverage, they either suffer from an older timespan, contain erroneous data, or have missing information. Massive-STEPS overcomes these issues by providing high-quality check-ins from 2012-2013 and 2017-2018, improving temporal quality for longitudinal POI recommendation studies. The dataset spans 12 diverse cities across multiple regions, with a focus on low-resource cities overlooked in previous research. Additionally, Massive-STEPS is enriched with metadata through alignment with Foursquare’s Open Source Places, providing crucial details such as POI geographical coordinates, POI names, and addresses. We also conducted an extensive benchmark on both supervised and zero-shot POI recommendation tasks, evaluating a wide range of models, including traditional methods, deep learning approaches, and recent LLM-based techniques. We further analyzed which urban features affect POI recommendation accuracy and found that cities with no dominant POI category tend to be harder to predict. By releasing this dataset and benchmark code publicly, we facilitate open and reproducible research, enabling future advancements in POI recommendation studies.
2 Related Works
Figure 1: Distribution of POI recommendation studies modeled on specific cities, filtered from Table IV in [68]. We identified and counted studies that explicitly mentioned city names, revealing the skewness of existing research, which is saturated around New York and Tokyo. In addition, we include the distribution of studies by LBSN platform, showing that Foursquare is by far the most commonly used source of check-in data. The list of identified studies is shown in Table 5.
2.1 Existing Datasets
@@
A survey conducted by [68] outlines the broad landscape of POI recommendation research, covering a wide range of models and architectures used in prior studies. While it offers a high-level overview of the datasets used, it lacks a dedicated discussion or evaluation of POI datasets. We address this gap by analyzing commonly used datasets in the field and positioning our dataset within this context.
@@
Building on the tabular summary provided in [68], which offers a representative overview of the broader POI recommendation literature, we investigated which datasets are most commonly used in prior studies. From their original table (Table IV), we filtered entries pertaining specifically to POI and next POI recommendation tasks and identified (1) the most frequently used LBSN check-in data sources and (2) the most commonly studied cities. As shown in Fig. 1, Foursquare remains the dominant source of LBSN data in existing studies, appearing in almost 50% of the surveyed works. While several variants of Foursquare datasets have been employed, the most widely used are the NYC and Tokyo Dataset [57] (often abbreviated as FSQ-NYC and FSQ-TKY) and the Global-scale Check-in Dataset (GSCD) [55, 56], curated by the same authors. Other LBSN sources occasionally used include Gowalla [4], Brightkite [4], and Weeplaces [26].
@@
Due to the widespread use of FSQ-NYC and FSQ-TKY [57], the majority of POI recommendation studies are disproportionately focused on these two cities, as illustrated in Fig. 1. While there is nothing inherently problematic about studying NYC and Tokyo, there has been growing interest in expanding research to a broader range of cities, particularly those that are underexplored or considered low-resource [63], as cultural and regional differences influence collective mobility behaviors. For instance, in some urban contexts, residents tend to commute to central business districts in the morning, whereas in others, nightlife activities such as visiting bars after work are more common [55]. Ensuring diverse geographic coverage is increasingly important, especially as LLMs are adopted for POI recommendation tasks. LLMs are known to exhibit geographical biases, often underperforming in regions with lower socioeconomic conditions [27]. It remains an open question whether LLM-based POI recommendation models can effectively generalize across diverse urban environments.
In addition, because many studies rely on the FSQ-NYC and FSQ-TKY, they are often constrained to the timespan it covers: check-in data from 2012 to 2013. However, POI data is inherently dynamic: venues may have closed, relocated, or changed in category over time. [61] underscores the importance of validating the temporal quality of POI datasets by recording whether and when a venue’s information has been updated to reflect real-world changes. This is particularly critical, as recommender systems should avoid suggesting POIs that no longer exist or have undergone substantial changes (e.g., a former bookstore converted into a coworking space) and behave dynamically over longitudinal periods [53]. Moreover, behavioral patterns captured over a decade ago may no longer align with modern user preferences and routines. For example, the opening of a new train station may significantly shift commuting patterns and the popularity of surrounding POIs.
@@
More recently, researchers have begun leveraging the broader Global-scale Check-in Dataset (GSCD) [55, 56], which spans 415 cities across 77 countries. Despite its wider geographic coverage, GSCD is temporally limited to the same 2012-2013 period as FSQ-NYC and FSQ-TKY, and thus suffers from similar issues of temporal quality. More critically, [29] demonstrated that GSCD suffers from significant data quality issues, with over 14 million check-ins (about 44%) of the dataset flagged as erroneous due to anomalous user behavior. These include (1) repeated check-ins at the same venue, (2) check-ins occurring within implausibly short time intervals (less than one minute), and (3) transitions between venues that would require travel speeds exceeding Mach 1, which are physically unreasonable.
To address these limitations, [29] introduced the Semantic Trails Dataset (STD), which applies systematic filtering procedures to enhance data quality. STD comprises two subsets: a cleaned version of GSCD covering 2012-2013 (STD 2013), and a newer collection of check-ins from 2017-2018 (STD 2018), sourced from Foursquare Swarm. STD 2018 also spans a wider range of cities, making it valuable for capturing globally distributed user behavior, in contrast to GSCD’s focus on densely populated urban centers. Both subsets follow the same rigorous filtering criteria, resulting in a higher-quality check-in dataset for POI recommendation tasks. Given these improvements, we adopted STD as the source for our check-in dataset.
@@
Another persistent challenge in POI recommendation research is the lack of reproducibility in dataset preprocessing. While some recent studies utilize datasets like GSCD to cover a wide range of cities, they often omit important details needed for replicating their data filtering processes. For example, AgentMove [7] conducted city-level filtering based on a minimum distance between each trajectory and the city center, but they did not specify how the city center was defined or what distance threshold was used. Generally, city filtering can be done in two ways: (1) by selecting points within a fixed radius from a central coordinate, or (2) by applying precise administrative boundaries, such as those from OpenStreetMap. The radius-based approach is less reliable, as it assumes a circular city shape and requires a subjective definition of the city’s center. In contrast, the boundary-based method is more robust and reproducible, and is the approach we adopted in this work. Without standardized, transparent filtering criteria and publicly available datasets, it remains difficult to ensure fair comparisons across POI recommendation methods.
2.2 Understanding Urban Features and POI Recommendation
POI recommendation studies that evaluate models across multiple city-level datasets often include analyses to assess how well their methods generalize across different urban contexts. It is well understood that POI recommendation accuracy metrics (e.g., Acc@k, NDCG@k) can vary substantially between cities and can be interpreted as a proxy for how easy or difficult a city is to model. The assumption is that higher performance reflects more predictable or structured mobility patterns. This viewpoint is consistent with prior work highlighting the role of cultural and urban-specific factors in shaping mobility behaviors [55, 34].
Several studies have proposed hypotheses connecting specific urban features to modeling difficulty. For example, GETNext [60] hypothesized that cities with fewer check-ins and higher spatial sparsity of POIs are harder to model. STHGCN [54] suggested that a larger number of user trajectories improves predictive accuracy by providing richer collaborative signals, whose architecture is designed to leverage. LLM4POI [22] proposed that cities with a greater variety of POI categories are easier to model due to LLMs’ contextual reasoning capabilities, whereas cities covering a broader geographic area tend to be more difficult to model. In the zero-shot POI recommendation setting, AgentMove  [7] reported two key findings: (1) geospatial biases inherent in LLMs can hinder prediction quality across cities, and (2) LLMs are influenced by city-specific mobility patterns.
Building on these insights, we used Massive-STEPS to explore how urban features affect POI recommendations. Its diverse set of 12 cities allows for a comprehensive analysis across different cultural and urban contexts. We analyzed the correlation between urban features and model accuracy, and based on the results, proposed a new hypothesis that contrasts previous findings in the literature.
3 Dataset
Table 2: Summary statistics of the 12 Massive-STEPS subsets, including the number of users, trajectories, POI locations, total check-ins, and train, validation, and test sample counts. For comparison, we also include statistics from existing Foursquare-based [57] and Gowalla-based [4] datasets. Due to variations in dataset preprocessing across studies, we report the version used in [54].
City Users Trajectories POIs Check-ins #train #val #test
NYC and Tokyo Check-in Dataset [57]
New York 1,048 14,130 4,981 103,941 72,206 1,400 1,347
Tokyo 2,282 65,499 7,833 405,000 274,597 6,868 7,038
Gowalla [4, 62]
California 3,957 45,123 9,690 238,369 154,253 3,529 2,780
Massive-STEPS
Beijing 56 573 1,127 1,470 400 58 115
Istanbul 23,700 216,411 53,812 544,471 151,487 21,641 43,283
Jakarta 8,336 137,396 76,116 412,100 96,176 13,740 27,480
Kuwait City 9,628 91,658 17,180 232,706 64,160 9,166 18,332
Melbourne 646 7,864 7,699 22,050 5,504 787 1,573
Moscow 3,993 39,485 17,822 105,620 27,639 3,949 7,897
New York 6,929 92,041 49,218 272,368 64,428 9,204 18,409
Petaling Jaya 14,308 180,410 60,158 506,430 126,287 18,041 36,082
São Paulo 5,822 89,689 38,377 256,824 62,782 8,969 17,938
Shanghai 296 3,636 4,462 10,491 2,544 364 728
Sydney 740 10,148 8,986 29,900 7,103 1,015 2,030
Tokyo 764 5,482 4,725 13,839 3,836 549 1,097
3.1 Creation Process

### v3

@@
Abstract.
Understanding human mobility through Point-of-Interest (POI) trajectory modeling is increasingly important for applications such as urban planning, personalized services, and generative agent simulation. However, progress in this field is hindered by two key challenges: the over-reliance on older datasets from 2012-2013 and the lack of reproducible, city-level check-in datasets that reflect diverse global regions. To address these gaps, we present Massive-STEPS (Massive Semantic Trajectories for Understanding POI Check-ins), a large-scale, publicly available benchmark dataset built upon the Semantic Trails dataset and enriched with semantic POI metadata. Massive-STEPS spans 15 geographically and culturally diverse cities and features more recent (2017-2018) and longer-duration (24 months) check-in data than prior datasets. We benchmarked a wide range of POI models on Massive-STEPS using both supervised and zero-shot approaches, and evaluated their performance across multiple urban contexts. By releasing Massive-STEPS, we aim to facilitate reproducible and equitable research in human mobility and POI trajectory modeling. The dataset and benchmarking code are available at: https://github.com/cruiseresearchgroup/Massive-STEPS.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
1. Introduction
@@
Human mobility data and POI trajectory modeling are essential for understanding how individuals interact with and move through physical spaces. This understanding enables a wide range of applications, including urban planning (Yuan et al., 2025), travel service recommendations (Feng et al., 2025), improved commercial advertising strategies (Yang et al., 2022b), and Point-of-Interest (POI) recommendation (Ding et al., 2020; Li et al., 2024; Zhang et al., 2025). Recently, human mobility data has become even more crucial with the increasing use of large language model (LLM) agents to simulate human-like behavior and routines (Zhou et al., 2024; Jiawei et al., 2024). However, while simulated and aggregated human mobility data are starting to gain popularity (Feng et al., 2020; Qin et al., 2023; Stanford et al., 2024; Jiang et al., 2025), they may not accurately reflect real-world, fine-grained individual human behavior (Salim et al., 2020), highlighting the value of evaluating on real-world data. These advancements are enabled by and large with Location-based Social Networks (LBSNs), which generate vast amounts of spatio-temporal data through user check-ins (Zhang et al., 2025; Li et al., 2024). This rich data source has allowed the development of POI recommendation systems that leverage users’ historical visiting behaviors to suggest relevant locations. Such systems enhance user engagement through personalization and provide commercial value to both users and businesses by aligning recommendations with individual preferences and available services (Ding et al., 2020).
@@
Our paper addresses three critical gaps in POI trajectory modeling research and datasets. First, as shown in Fig. 7, the field remains dominated by studies focused on just two cities, New York and Tokyo, based on the Foursquare dataset curated by Yang et al. (2014). This dataset, collected in 2012-2013, raises concerns about its temporal quality, as many POIs may no longer exist and user behavior may have changed (Yeow et al., 2021). While some recent studies have expanded to other cities (Zhang et al., 2024a; Merinov and Ricci, 2024; Feng et al., 2025), they often rely on the Global-scale Check-in Dataset (GSCD) (Yang et al., 2015, 2016), which, despite its extensive coverage, is also from 2012-2013 and contains nearly 50% erroneous entries (Monti et al., 2018). Second, most existing studies are difficult to reproduce, either due to the lack of clearly defined geographic boundaries or the unavailability of the datasets themselves, hindering fair comparison and replication. Finally, we join recent efforts (Yuan et al., 2025) in advocating for the inclusion of low-resource and underrepresented cities. Expanding beyond well-studied urban centers is essential for building more generalizable and universally applicable POI models. Table 1 summarizes these limitations in terms of geographic coverage, temporal span, and reproducibility.
Figure 1. Massive-STEPS Dataset and Benchmark Overview.
Massive-STEPS Dataset and Benchmark Overview.
Massive-STEPS Dataset
In this paper, we introduced the Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) Dataset, derived from the Semantic Trails dataset (STD) (Monti et al., 2018). Massive-STEPS includes high-quality check-ins from 2012-2013 and 2017-2018, providing more modern and updated POI check-in data. This supports longitudinal POI trajectory modeling studies and addresses the limitations of older datasets commonly used in prior studies. The dataset covers 15 diverse cities across multiple regions, including East, West, and Southeast Asia, North and South America, Australia, the Middle East, and Europe. Notably, we placed a deliberate emphasis on under-explored regions by including cities such as Jakarta, Kuwait City, and Petaling Jaya, filling a key gap in POI trajectory research that has largely focused on major urban centers. We further enriched STD by aligning it with Foursquare’s Open Source Places dataset, incorporating missing metadata such as geographic coordinates, POI names, and addresses.
Benchmark Tasks
To demonstrate the utility of this dataset, we conducted an extensive benchmark on three tasks as shown in Fig. 1: (1) supervised POI recommendation, (2) zero-shot POI recommendation, and (3) spatiotemporal classification and reasoning. Our benchmark covers a wide range of models, including traditional approaches, deep learning-based models, and more recent LLM-based methods. The goal of POI recommendation task is to predict a set of POIs that a user is likely to visit based on their current check-in trajectory and historical behavior. This reflects real-world applications such as personalized POI recommendations in location-based services. Similarly, the goal of spatiotemporal classification and reasoning is to assess how effectively models (e.g., LLMs) leverage, interpret, and reason about POI trajectories. In addition, the scale of our dataset allows us to examine how urban features influence POI modeling accuracy. Building on prior hypotheses, we propose a new insight: cities with more evenly distributed POI categories tend to be harder to model, as the absence of a dominant POI category makes user behavior less predictable.
Contribution
This paper introduces the Massive Semantic Trajectories for Understanding POI Check-ins (Massive-STEPS) dataset, addressing gaps in existing POI trajectory modeling research. Current POI check-in datasets are often only from 2012-2013, skewed to a few cities, and lack semantic metadata, hindering the development of robust and globally applicable models. While datasets like GSCD and STD offer broad geographic coverage, they either suffer from an older timespan, contain erroneous data, or have missing information. Massive-STEPS overcomes these issues by providing high-quality check-ins from 2012-2013 and 2017-2018, improving temporal quality for longitudinal POI trajectory modeling studies. The dataset spans 15 diverse cities across multiple regions, with a focus on low-resource cities overlooked in previous research. Additionally, Massive-STEPS is enriched with metadata through alignment with Foursquare’s Open Source Places, providing crucial details such as POI geographical coordinates, POI names, and addresses. We also conducted an extensive benchmark on both supervised and zero-shot POI recommendation and trajectory classification tasks, evaluating a wide range of models, including traditional methods, deep learning approaches, and recent LLM-based techniques. We further analyzed which urban features affect POI modeling accuracy and found that cities with no dominant POI category tend to be harder to predict. By releasing this dataset and benchmark code publicly, we facilitate open and reproducible research, enabling future advancements in urban mobility studies.
Table 1. Comparison of check-in datasets commonly used for POI modeling tasks.
@@
#cities Years #months POI Attributes Replicable Open
NYC/TKY (Yang et al., 2014) 2 2012-2013 11   Coordinates, Category
Gowalla-CA (Yuan et al., 2013) 1 2009-2010 21   Coordinates, Category
AgentMove (Feng et al., 2025) 12 2012-2013 17   Coordinates, Category
Massive-STEPS 15 2012-2013, 2017-2018 24 Coordinates, Category, Name, Address
2. Existing Datasets
@@
A survey conducted by (Zhang et al., 2025) outlines the landscape of POI trajectory modeling research, covering a wide range of models and architectures used in prior studies. While it offers a high-level overview of the datasets used, it lacks a dedicated discussion or evaluation of POI datasets. We address this gap by analyzing commonly used datasets and positioning our dataset within this context.
Table 2. Overview of POI dynamics: total POIs ever opened (according to Foursquare OS Places), POIs confirmed closed by 2025, and POIs closed during 2014-2016, corresponding to the temporal gap in our dataset.
City POIs Opened Total POIs Closed Closed within 2014-2016
New York 49,218 13,009 (26.43%) 3,118 (6.34%)
Melbourne 7,699 1,850 (24.03%) 209 (2.71%)
Sydney 8,986 1,759 (19.57%) 253 (2.82%)
Moscow 17,822 3,021 (16.95%) 868 (4.87%)
São Paulo 38,377 4,990 (13.00%) 1,257 (3.28%)
Shanghai 4,462 661 (14.81%) 81 (1.82%)
Tokyo 4,725 421 (8.91%) 0 (0.00%)
Petaling Jaya 60,158 4,186 (6.96%) 1,533 (2.55%)
Istanbul 53,812 2,833 (5.26%) 481 (0.89%)
Beijing 1,127 56 (4.97%) 10 (0.89%)
Jakarta 76,116 3,527 (4.63%) 483 (0.63%)
Bandung 29,026 1,053 (3.63%) 182 (0.63%)
Palembang 4,343 143 (3.29%) 23 (0.53%)
Tangerang 12,956 383 (2.96%) 50 (0.39%)
Kuwait City 17,180 161 (0.94%) 22 (0.13%)
@@
Building on the table from (Zhang et al., 2025), we investigated which datasets are most commonly used in prior studies. From their original table (Table IV), we filtered entries pertaining specifically to POI and next POI recommendation tasks and identified (1) the most frequently used LBSN check-in data sources and (2) the most commonly studied cities. As shown in Fig. 7, Foursquare remains the dominant source of LBSN data in existing studies, appearing in almost 50% of the surveyed works. While several variants of Foursquare datasets have been employed, the most widely used are the NYC and Tokyo Dataset (Yang et al., 2014) (often abbreviated as FSQ-NYC and FSQ-TKY) and the Global-scale Check-in Dataset (GSCD) (Yang et al., 2015, 2016), curated by the same authors. Other LBSN sources occasionally used include Gowalla (Cho et al., 2011), Brightkite (Cho et al., 2011), and Weeplaces (Liu et al., 2017).
@@
Due to the widespread use of FSQ-NYC and FSQ-TKY (Yang et al., 2014), the majority of POI trajectory studies are disproportionately focused on these two cities, as illustrated in Fig. 7. While there is nothing inherently problematic about NYC and Tokyo, there has been growing interest in expanding research to a broader range of cities, particularly those that are underexplored or considered low-resource (Yuan et al., 2025), as cultural and regional differences influence collective mobility behaviors. For instance, in some cities, residents tend to commute to business districts in the morning, whereas in others, nightlife activities such as visiting bars after work are more common (Yang et al., 2015). Ensuring diverse geographic coverage is increasingly important, especially as LLMs are adopted for POI trajectory modeling tasks. LLMs are known to exhibit geographical biases against regions with lower socioeconomic conditions (Manvi et al., 2024). Whether LLMs can generalize across diverse urban environments is to be investigated.
In addition, because many studies rely on the FSQ-NYC and FSQ-TKY, they are often constrained to the timespan it covers: check-in data from 2012 to 2013. However, POI data is inherently dynamic: venues may have closed, relocated, or changed in category over time. Table 2 illustrates this by detailing the total number of POIs ever opened, the number of POIs closed by 2025, and the number of POIs that closed between 2014 and 2016 alone. These substantial closure rates in major cities highlight the critical need for newer, multi-period datasets that reflect the actual lifecycle of POIs.
Yeow et al. (2021) further underscores the importance of validating the temporal quality of POI datasets by recording whether and when a venue’s information has been updated to reflect real-world changes. This is particularly critical, as recommender systems should avoid suggesting POIs that no longer exist or have undergone substantial changes (e.g., a former bookstore converted into a coworking space) and behave dynamically over longitudinal periods (Yabe et al., 2024). Moreover, visitation patterns may change over time depending on the city (see Fig.2), potentially due to shifts in routines. For instance, the opening of a new train station can noticeably alter commuting patterns and the popularity of nearby POIs.
Figure 2. Top 10 most visited POI categories in Bandung and Kuwait City across two time periods. Remaining cities are shown in Appendix B.
Top 10 most visited POI categories in Bandung and Kuwait City across two time periods.
@@
More recently, researchers have begun leveraging the broader Global-scale Check-in Dataset (GSCD) (Yang et al., 2015, 2016), which spans 415 cities across 77 countries. Despite its wider geographic coverage, GSCD is temporally limited to the same 2012-2013 period as FSQ-NYC and FSQ-TKY, and thus suffers from similar issues of temporal quality. More critically, Monti et al. (2018) demonstrated that GSCD suffers from significant data quality issues, with over 14 million check-ins (about 44%) of the dataset flagged as erroneous due to anomalous user behavior. These include (1) repeated check-ins at the same venue, (2) check-ins occurring within implausibly short time intervals (less than one minute), and (3) transitions between venues that would require travel speeds exceeding Mach 1, which are physically impossible.
To address these limitations, Monti et al. (2018) introduced the Semantic Trails Dataset (STD), which applies systematic filtering procedures to enhance data quality. STD comprises two subsets: a cleaned version of GSCD covering 2012-2013 (STD 2013), and a newer collection of check-ins from 2017-2018 (STD 2018), sourced from Foursquare Swarm. STD 2018 also spans a wider range of cities, making it valuable for capturing globally distributed user behavior, in contrast to GSCD’s focus on densely populated urban centers. Both subsets follow the same rigorous filtering criteria, resulting in a higher-quality check-in dataset for downstream POI trajectory modeling tasks. Given these improvements, we adopted STD as the source for our check-in dataset.
@@
Another persistent challenge in POI trajectory research is the lack of reproducibility in dataset preprocessing. While some recent studies utilize datasets like GSCD to cover a wide range of cities, they often omit important details needed for replicating their data filtering processes. For example, (Feng et al., 2025) and (Zuo and Zhang, 2024) conducted city-level filtering, but they did not specify how the city boundaries were defined or what distance-based thresholds were used. Similarly, the Weeplaces dataset used in (Chen and Zhu, 2025) and (Cao et al., 2023) is no longer available. To further support this claim, we provide an extensive list of dataset reproducibility issues in all the studies reviewed in (Zhang et al., 2025), in Table 8. As shown, almost none of the datasets used in these works are fully reproducible or publicly available, except for FSQ-NYC/TKY (Yang et al., 2014) and Gowalla-CA (Yuan et al., 2013), leading to a heavy reliance on these datasets.
3. Massive-STEPS Dataset
3.1. Creation Process